OpenStudy (anonymous):

Help any one I need help solving the system of equation. x+y+z=11 y+z=10 z=6

OpenStudy (anonymous):

$$x=1$$ $$z=6$$ $$y=4$$

OpenStudy (anonymous):

Use Gaussian elimination

OpenStudy (anonymous):

Hmm, i'm not sure what gaussiam elimination is but thank you!!!

OpenStudy (nurali):

x+y+z=11---------(1) y+z=10-----------(2) z=6--------------(3) put z=6 in equation (2) y+6=10 subtract both sides by '6' y+6-6=10-6 y=4 put y and z in equation '1' x+4+6=11 x+10=11 x=11-10 x=1